# Cosmologically Coupled Black Holes with Regular Horizons

### v1

@@ I Introduction
It is an intriguing possibility that black holes (BHs) and other compact astrophysical objects may couple to the large-scale Friedmann-Lemaître-Robertson-Walker (FLRW) cosmological dynamics [McVittie:1933zz, Einstein:1945id, Einstein:1946zz, Croker:2024jfg, Faraoni:2023hin, Calza:2024qxn, Calza:2024xdh, Calza:2024fzo, dicke1964evolution, Vaidya:1968zza, DEath:1975jps, Gautreau:1984pny, Cooperstock:1998ny, Nayak:2000mr, Baker:2000yh, Bolen:2000dz, Dominguez:2001it, Ellis:2001cq, Gao:2004cr, Sheehan:2004wa, Nesseris:2004uj, Sultana:2005tp, Li:2006zh, Adkins:2006kw, McClure:2006kg, Sereno:2007tt, Faraoni:2007es, Balaguera-Antolinez:2007csw, Mashhoon:2007qm, Carrera:2008pi, Gao:2011tq, Faraoni:2014nba, Kopeikin:2014qna, Faraoni:2015saa, Mello:2016irl, Faraoni:2018xwo, Guariento:2019ock, Spengler:2021vxy, Agatsuma:2022ewd, Croker:2019mup, Croker:2020, Croker:2020plg, Farrah:2023opk, Wang:2023aqe, Gaur:2023hmk, Parnovsky:2023wkc, Avelino:2023rac, Dahal:2023hzo, Gao:2023keg, Cadoni:2023lum, Cadoni:2023lqe, Cadoni:2024jxy, Cadoni:2024rri]. This idea, originally introduced by McVittie [McVittie:1933zz], has recently regained attention within a genuinely dynamical setting. The cosmological embedding (CE) of BHs is usually constructed starting from a static, spherically-symmetric model and employing a set of physically motivated assumptions. It has been proposed that the mass $M$ of these cosmologically coupled (CC) astrophysical objects could vary with the redshift [Croker:2019mup, Cadoni:2023lum, Cadoni:2023lqe], independently of accretion processes. This opened a new window to test fundamental properties of BHs and their mimickers, as well as the evolution of their mass spectrum at different redshifts [Cadoni:2023lum].
General arguments based on causality and an effective fluid description suggest a power-law scaling of the mass with the scale factor $a$, $M\sim a^{k}$ with $3\leq k\leq 3$ [Croker:2019mup]. These theoretical predictions have been tested against observations using both electromagnetic and gravitational-wave channels [Farrah:2023opk, Cadoni:2023lum, Rodriguez:2023gaa, Andrae:2023wge, Lei:2023mke, Amendola:2023ays, Lacy:2023kbb, Calza:2024qxn], yielding, however, no conclusive evidence yet. The main issue at the observational level is the intrinsic difficulty in comparing homogeneous samples of BH mass spectra at different redshifts.
@@
In the present work, we show that the theoretical difficulty mentioned above can be overcome. Inspired by the approach of [deLimaRodrigues2025, Cadoni:2023lum], we construct solutions describing spherical objects embedded in a cosmological background with no radial matter flow (i.e., no accretion). We derive a general parametrization of the field equations in terms of a local mass function, together with the density and pressure components of the cosmological fluid. This parametrization is particularly useful being directly tied to physical intuition, as it employs the standard cosmological time coordinate, the comoving radius and a natural definition of the mass of the CE object.
Our construction is based on a given regular Misner-Sharp mass [Misner:1964je], and is shown here to generate solutions that are regular everywhere, apart from the possible central singularity. As a consequence, when applied to BHs, their event horizons remain regular despite the embedding, thus solving the previously mentioned issue of CC BH models. We discuss both BHs with a central singularity and nonsingular BHs. Finally, we derive a new CE of the Schwarzschild BH, distinct from the well-known McVittie solution [McVittie:1933zz].
@@ II Spacetime and the energy momentum tensor
Here, $\alpha(\eta,r)$ and $\beta(\eta,r)$ are metric functions depending on the conformal time $\eta$ and $r$, while $a(\eta)$ is the scale factor. The source of the gravitational field is an anisotropic fluid, which is described by the energy-momentum tensor
@@
where $u_{\mu}$ and $w_{\mu}$ satisfy the normalization conditions $g^{\mu\nu}u_{\mu}u_{\nu}=-1$, $g^{\mu\nu}w_{\mu}w_{\nu}=1$, and $u^{\mu}w_{\mu}=0$. Here, $\rho$, $p_{\parallel}$, and $p_{\perp}$ denote the energy density and the radial and transverse pressures of the fluid, respectively.
The Einstein field equations following from the metric ansatz (1) sourced by the energy–momentum tensor (2) can be written as [Cadoni:2020jxe] (dot and prime will refer to derivation with respect to $\eta$ and $r$, respectively)
@@ II.1 Parametrization of the field equations in terms of local mass functions
After some manipulation (explicit details can be found in Appendix A), the combination of Einstein’s equation and covariant conservation of energy-momentum tensor leads to
@@
The system of equations admits a uniquely-defined solution once the free function $M_{\mbox{\tiny MS}}^{l}$ is specified, together with suitable boundary conditions. In particular, in order to retrieve FLRW cosmology at large radial distances, we must necessarily consider models satisfying $\alpha,\,\beta\to 0$ as $r\to\infty$.
@@
which is defined in terms of the usual Misner-Sharp (MS) mass [Misner:1964je]
@@ II.2 Exact solutions
Let us conclude the section with some comments on the physical properties of $\mathcal{M}(r,a)$. Although having been initially introduced merely as a tool to simplify the integration of the field equations, it acquires a clear physical meaning when evaluated at the apparent horizon. For spherically-symmetric metrics, the location of the latter is determined by the condition [Faraoni_2009]
$\nabla_{\mu}\ell\,\nabla^{\mu}\ell$ $=1-\frac{2M_{\mbox{\tiny MS}}(r_{\text{h}},a)}{\ell}=$
@@ III Cosmologically embedded black holes
Let us now focus on the subclass of solutions describing the CE of BHs, including both the singular Schwarzschild and nonsingular BH models, like those analyzed in Ref. [Cadoni:2022chn]. The simplest way to contruct CE of BH solutions is to start from a mass profile $M(r)$ describing a static, spherically-symmetric BH and promote the radial coordinate $r\to ra=\ell$. This method has been firstly introduced by McVittie, working however in isotropic coordinates [McVittie:1933zz]. Notice that, since after the embedding, $\alpha$ and $\beta$ will be functions of $\ell$ only, the condition $\alpha=-\beta$, immediately follows from Eq. 3a, ensuring the absence of radial energy flux. In this way, we can construct two different classes of exact solutions. Despite the fact that the first one can be obtained as a particular case of Eq. 20, it is easier to derive it directly from Einstein’s field equations
@@
A first physically relevant case is obtained by choosing $M(\ell)=M_{0}=\text{const.}$, which corresponds to the CE of the Schwarzschild BH. However, the resulting spacetime differs from the McVittie solution [McVittie:1933zz], as the two cannot be related by any coordinate transformation. This inequivalence arises because the McVittie spacetime is derived under the assumption of vanishing energy flux in the isotropic radial coordinate and is sourced by an isotropic fluid. In contrast, our solutions enforce the no-flux condition in the comoving areal coordinate and are sourced by an anisotropic fluid. When expressed in terms of the proper length $\ell$, our solution reads
@@
with $t$ the cosmological time. Interestingly, it bears some similarities with the Sultana-Dyer solution [SultanaDyer2005, Faraoni_2009], which, however, describes the cosmological embedding of a BH with non-zero energy fluxes.
@@
This solution is now free of curvature singularities, and outside the apparent horizon we have $\partial_{a}\gamma(\ell,a)\sim\partial_{a}\rho_{\text{c}}\approx 0$, so that the contribution coming from the integral in Eq. 28c becomes negligible. This is a quite important result and confirms the findings above: the horizon singularity - observed in all previously studied BH embeddings - is resolved once one considers solutions that consistently incorporate the dynamical interplay between the local metric and the large-scale cosmological dynamics. Therefore, the curvature singularity at the horizon, featured by the McVittie solution (see, e.g., Ref. [Modesto:2025cre]), is here cured.
@@
The results presented so far naturally extend to the CE of nonsingular BHs with de Sitter core (see, e.g., Ref. [Cadoni:2023lum]). They are characterized by a mass function $M(r)$ behaving as $r^{3}$ near $r=0$ and as $M_{0}$ as $r\to\infty$. Although a possible CE for such models has already been proposed in literature [Cadoni:2023lum, Cadoni:2023lqe, Cadoni:2024rri], these embeddings still exhibit a curvature singularity at the horizon  [^1] This feature has not been previously emphasized, but it can be readily verified by computing the curvature invariants of the solutions presented in Ref. [Cadoni:2023lqe].. Remarkably, the embedding introduced in the present work removes this pathology and the horizon singularities appearing in the CE of otherwise regular BH models are eliminated.
Finally, our results provide further insight into the theorem of Ref. [Faraoni:2024ghi], which forbids the existence of regular static horizons in any CE of black hole solution. The present construction is fully consistent with this theorem, since the solutions obtained here do not admit static event horizons, but only apparent horizons. At the same time, our results explain the singular behaviour at the horizon of the solution (25) as a direct consequence of neglecting the, $\rho_{\text{c}}$-dependent contribution.
@@ IV Energy conditions
We first consider the CC solutions (20) and assume that terms proportional to derivatives of $M_{\mbox{\tiny MS}}^{l}$ can be neglected. This assumption is justified, for instance, when $M_{\mbox{\tiny MS}}^{l}=M_{0}=\text{const.}$, and more generally holds for models in which the density profile decays sufficiently rapidly outside the BH. We get
@@
Let us now turn to the class of solutions (25). In this case one finds $e^{\alpha}(p_{\parallel}+\rho)=p_{\text{c}}+\rho_{\text{c}}.$ The NEC is, therefore, always satisfied in the BH exterior. On the other hand, $\rho+p_{\perp}$ depends on $k$ and its $r$-derivative. In general, it will always exist a value $\ell_{0}$, typically of the order of $\ell_{\text{h}}$ such that for $\ell\leq\ell_{0}$, the inequality $\rho+p_{\perp}\geq 0$ is violated. Thus, we expect the NEC (hence the WEC) to be violated in the near-horizon region. A similar argument shows that this is also true for the DEC.
@@ V Conclusions
Our results also elucidate the physical content of the theorem proved in Ref. [Faraoni:2024ghi], which forbids the existence of regular static horizons in CE solutions. On the one hand, our general CE solution fully complies with the theorem, as it admits an apparent, but not a static horizon. On the other hand, our analysis clarifies the fate of the event horizon of a local static BH under cosmological embedding. Prior to the embedding, the BH event horizon, defined by the condition $2M(r_{\text{H}})=r_{\text{H}}$, is a completely regular surface. If the CE is performed while neglecting backreaction, namely by employing Eq. 25, the resulting spacetime develops a curvature singularity at the surface defined by $2M(\ell_{\text{h}})=\ell_{\text{h}}$. This singular behavior is avoided once backreaction effects are properly included, that is, when the CE is constructed using Eq. 28.
@@ Acknowledgements.
Appendix A The Misner-Sharp mass parametrization

### v2

@@ I Introduction
It is an intriguing possibility that black holes (BHs) and other compact astrophysical objects may couple to the large-scale Friedmann-Lemaître-Robertson-Walker (FLRW) cosmological dynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. This idea, originally introduced by McVittie [1], has recently regained attention within a genuinely dynamical setting. The cosmological embedding (CE) of BHs is usually constructed starting from a static, spherically-symmetric model and employing a set of physically motivated assumptions. It has been proposed that the mass $M$ of these cosmologically coupled (CC) astrophysical objects could vary with the redshift [40, 50, 51], independently of accretion processes. This opened a new window to test fundamental properties of BHs and their mimickers, as well as the evolution of their mass spectrum at different redshifts [50].
General arguments based on causality and an effective fluid description suggest a power-law scaling of the mass with the scale factor $a$, $M\sim a^{k}$ with $3\leq k\leq 3$ [40]. These theoretical predictions have been tested against observations using both electromagnetic and gravitational-wave channels [43, 50, 56, 57, 58, 59, 60, 6], yielding, however, no conclusive evidence yet. The main issue at the observational level is the intrinsic difficulty in comparing homogeneous samples of BH mass spectra at different redshifts.
@@
In the present work, we show that the theoretical difficulty mentioned above can be overcome. Inspired by the approach of [61, 50], we construct solutions describing spherical objects embedded in a cosmological background with no radial matter flow (i.e., no accretion). We derive a general parametrization of the field equations in terms of a local mass function, together with the density and pressure components of the cosmological fluid. This parametrization is particularly useful being directly tied to physical intuition, as it employs the standard cosmological time coordinate, the comoving radius and a natural definition of the mass of the CE object.
Our construction is based on a given regular Misner-Sharp mass [62], and is shown here to generate solutions that are regular everywhere, apart from the possible central singularity. As a consequence, when applied to BHs, their event horizons remain regular despite the embedding, thus solving the previously mentioned issue of CC BH models. We discuss both BHs with a central singularity and nonsingular BHs. Finally, we derive a new CE of the Schwarzschild BH, distinct from the well-known McVittie solution [1].
@@ II Spacetime and the energy momentum tensor
Here, $\alpha(\eta,r)$ and $\beta(\eta,r)$ are metric functions depending on the conformal time $\eta$ and $r$, while $a(\eta)$ is the scale factor. The source of the gravitational field is an anisotropic fluid, which is described by the energy-momentum tensor (e.g., [63, 64, 65] and references therein)
@@
where $u_{\mu}$ and $w_{\mu}$ satisfy the normalization conditions $g^{\mu\nu}u_{\mu}u_{\nu}=-1$, $g^{\mu\nu}w_{\mu}w_{\nu}=1$, and $u^{\mu}w_{\mu}=0$. Here, $\rho$, $p_{\parallel}$, and $p_{\perp}$ denote the energy density and the radial and transverse pressures of the fluid, respectively. All the three scalar quantities $\rho$, $p_{\parallel}$, and $p_{\perp}$, are at this point independent. The above $T_{\mu\nu}$ is the most general spherically-symmetric energy-momentum tensor without energy flux. As we will see below, $\rho,\,p_{\parallel},\,p_{\perp}$ are not dynamically independent, but they are constrained by covariant conservation of the stress-energy tensor (see Eqs. 3d and 3e below). Notice also that anisotropy of the source in Einstein’s equation is achieved by taking $p_{\perp}\neq p_{\parallel}$, which generates an anisotropic stress, while the spacetime metric (1) remains isotropic.
The Einstein field equations following from the metric ansatz (1) sourced by the energy–momentum tensor (2) can be written as [66] (see also Appendix A)
@@
In the above, dot and prime refer to derivation with respect to $\eta$ and $r$, respectively. The above equations are valid for any value of the spatial curvature parameter $\mathcal{K}$ (i.e., $\mathcal{K}=0$ or $\mathcal{K}=\pm 1$). For $\mathcal{K}\not=0$, $\rho$, $p_{\parallel}$ and $p_{\perp}$, depend on a geometric-contribution term that behaves like $\mathcal{K}a^{-2}$.
@@ II.1 Parametrization of the field equations in terms of local mass functions
In the above equations, the additional contribution from a possibly non-zero spatial curvature ($\mathcal{K}=\pm 1$) is included in $\rho_{c}$ and $p_{c}$.
After some manipulation (explicit details can be found in Appendix B), the combination of Einstein’s equation and covariant conservation of energy-momentum tensor leads to
@@
The system of equations admits a uniquely-defined solution once the free function $M_{\mbox{\tiny MS}}^{l}$ is specified, together with suitable boundary conditions. In particular, in order to retrieve spatially-flat ($\mathcal{K}=0$) FLRW cosmology at large radial distances, we must necessarily consider models satisfying $\alpha,\,\beta\to 0$ as $r\to\infty$. For FLRW cosmologies with non-vanishing spatial curvature ($\mathcal{K}=\pm 1$), one imposes that $\alpha\to 0$ and $e^{\beta}\to 1/(1-\mathcal{K}r^{2})$ for sufficiently large $r$. In such limit, $\frac{2M_{\mbox{\tiny MS}}^{l}}{ra}$ must reduce to $\mathcal{K}r^{2}$. From Eq. (6), one sees that $\rho\to\rho_{c}+\frac{3\mathcal{K}}{8\pi a^{2}}$, $p_{\parallel}\to p_{c}-\frac{\mathcal{K}}{8\pi a^{2}}$ and $p_{\perp}\to p_{c}-\frac{\mathcal{K}}{8\pi a^{2}}$ as long as $\alpha\to 0$ and $e^{\beta}\to(1-\mathcal{K}r^{2})^{-1}$. In the following sections, we shall restrict our analysis to the case $\mathcal{K}=0$, as this is the most relevant case for the late universe [67].
@@
which is defined in terms of the usual Misner-Sharp (MS) mass [62]
@@ II.2 Exact solutions
Let us conclude the section with some comments on the physical properties of $\mathcal{M}(r,a)$. Although having been initially introduced merely as a tool to simplify the integration of the field equations, it acquires a clear physical meaning when evaluated at the apparent horizon. For spherically-symmetric metrics, the location of the latter is determined by the condition [68, 69]
$\nabla_{\mu}\ell\,\nabla^{\mu}\ell$ $=1-\frac{2M_{\mbox{\tiny MS}}(r_{\text{h}},a)}{\ell}$
@@ III Cosmologically embedded black holes
Let us now focus on the subclass of solutions describing the CE of BHs, including both the singular Schwarzschild and nonsingular BH models, like those analyzed in Ref. [70]. The simplest way to contruct CE of BH solutions is to start from a mass profile $M(r)$ describing a static, spherically-symmetric BH and promote the radial coordinate $r\to ra=\ell$. This method has been firstly introduced by McVittie, working however in isotropic coordinates [1]. Notice that, since after the embedding, $\alpha$ and $\beta$ will be functions of $\ell$ only, the condition $\alpha=-\beta$, immediately follows from Eq. 3a, ensuring the absence of radial energy flux. In this way, we can construct two different classes of exact solutions. Despite the fact that the first one can be obtained as a particular case of Eq. 20, it is easier to derive it directly from Einstein’s field equations
@@
A first physically relevant case is obtained by choosing $M(\ell)=M_{0}=\text{const.}$, which corresponds to the CE of the Schwarzschild BH. However, the resulting spacetime differs from the McVittie solution [1], as the two cannot be related by any coordinate transformation. This inequivalence arises because the McVittie spacetime is derived under the assumption of vanishing energy flux in the isotropic radial coordinate and is sourced by an isotropic fluid. In contrast, our solutions enforce the no-flux condition in the comoving areal coordinate and are sourced by an anisotropic fluid. When expressed in terms of the proper length $\ell$, our solution reads
@@
with $t$ the cosmological time. Interestingly, it bears some similarities with the Sultana-Dyer solution [71, 68], which, however, describes the cosmological embedding of a BH with non-zero energy fluxes.
@@
This solution is now free of curvature singularities, and outside the apparent horizon we have $\partial_{a}\gamma(\ell,a)\sim\partial_{a}\rho_{\text{c}}\approx 0$, so that the contribution coming from the integral in Eq. 28c becomes negligible. This is a quite important result and confirms the findings above: the horizon singularity - observed in all previously studied BH embeddings - is resolved once one considers solutions that consistently incorporate the dynamical interplay between the local metric and the large-scale cosmological dynamics. Therefore, the curvature singularity at the horizon, featured by the McVittie solution (see, e.g., Ref. [72]), is here cured.
@@
The results presented so far naturally extend to the CE of nonsingular BHs with de Sitter core (see, e.g., Ref. [50]). They are characterized by a mass function $M(r)$ behaving as $r^{3}$ near $r=0$ and as $M_{0}$ as $r\to\infty$. Although a possible CE for such models has already been proposed in literature [50, 51, 53], these embeddings still exhibit a curvature singularity at the horizon [^1] This feature has not been previously emphasized, but it can be readily verified by computing the curvature invariants of the solutions presented in Ref. [51].. Remarkably, the embedding introduced in the present work removes this pathology and the horizon singularities appearing in the CE of otherwise regular BH models are eliminated.
Finally, our results provide further insight into the theorem of Ref. [73], which forbids the existence of regular static horizons in any CE of black hole solution. The present construction is fully consistent with this theorem, since the solutions obtained here do not admit static event horizons, but only apparent horizons. At the same time, our results explain the singular behaviour at the horizon of the solution (25) as a direct consequence of neglecting the, $\rho_{\text{c}}$-dependent contribution.
@@ IV Energy conditions
We first consider the CC solutions (20) and assume that terms proportional to derivatives of $M_{\mbox{\tiny MS}}^{l}$ can be neglected. This assumption is justified, for instance, when $M_{\mbox{\tiny MS}}^{l}=M_{0}=\text{const.}$, and more generally holds for models in which the density profile decays sufficiently rapidly outside the BH. We get, using the expressions for $\rho,\,p_{\parallel},\,p_{\perp}$ given by (6),
@@
Let us now turn to the class of solutions (25). In this case one finds using the expressions for $\rho,\,p_{\parallel},\,p_{\perp}$ given by Eq. 25, $e^{\alpha}(p_{\parallel}+\rho)=p_{\text{c}}+\rho_{\text{c}}.$ The NEC is, therefore, always satisfied in the BH exterior. On the other hand, $\rho+p_{\perp}$ depends on $k$ and its $r$-derivative. In general, it will always exist a value $\ell_{0}$, typically of the order of $\ell_{\text{h}}$ such that for $\ell\leq\ell_{0}$, the inequality $\rho+p_{\perp}\geq 0$ is violated. Thus, we expect the NEC (hence the WEC) to be violated in the near-horizon region. A similar argument shows that this is also true for the DEC.
@@ V Conclusions
Our results also elucidate the physical content of the theorem proved in Ref. [73], which forbids the existence of regular static horizons in CE solutions. On the one hand, our general CE solution fully complies with the theorem, as it admits an apparent, but not a static horizon. On the other hand, our analysis clarifies the fate of the event horizon of a local static BH under cosmological embedding. Prior to the embedding, the BH event horizon, defined by the condition $2M(r_{\text{H}})=r_{\text{H}}$, is a completely regular surface. If the CE is performed while neglecting backreaction, namely by employing Eq. 25, the resulting spacetime develops a curvature singularity at the surface defined by $2M(\ell_{\text{h}})=\ell_{\text{h}}$. This singular behavior is avoided once backreaction effects are properly included, that is, when the CE is constructed using Eq. 28.
This work is not the first to investigate cosmological coupling and use a mass-related function to parametrize possible GR solutions, as we did in Eqs. (6, 20) and as previously considered in Ref. [74]. However, the analysis in Ref. [74] assumes isotropic pressure, which prevents the avoidance of singularities beyond the central one for $\mathcal{K}=0$ [75, 76]. In contrast, we have developed a new parametrization of GR solutions that relaxes the assumption of isotropic pressure. We have shown that this generalized framework can eliminate all singularities—apart from the central one—even in the case $\mathcal{K}=0$. Consequently, our results provide a partial extension of the $\mathcal{K}=0$ solutions of Ref. [74] to the case of an anisotropic fluid.
@@ Acknowledgements.
Appendix A Detailed derivation of the field equations
The field equations (3) were previously derived using similar notation and assumptions in Refs. [66, 50, 52]. For completeness and clarity, we present the complete derivation here. A Mathematica notebook is also provided  [^2] https://github.com/davi-rodrigues/ccbh-rh. which allows to reproduce the following results. For brevity, we use the index notation $(t,r,\theta,\phi)\leftrightarrow(0,1,2,3)$.
The non-vanishing Christoffel symbols read
$\Gamma^{0}_{00}$ $=\frac{\dot{a}}{a}+\frac{1}{2}\dot{\alpha},\qquad\Gamma^{0}_{01}=\frac{1}{2}\alpha^{\prime},$ (31a)
$\Gamma^{0}_{11}$ $=\frac{e^{-\alpha+\beta}}{2a}\left(2\dot{a}+a\dot{\beta}\right),$ (31b)
$\Gamma^{0}_{22}$ $=\frac{e^{-\alpha}r^{2}\dot{a}}{a},\qquad\Gamma^{0}_{33}=\frac{e^{-\alpha}r^{2}\sin^{2}\theta\,\dot{a}}{a},$ (31c)
$\Gamma^{1}_{01}$ $=\frac{\dot{a}}{a}+\frac{1}{2}\dot{\beta},\qquad\Gamma^{2}_{02}=\Gamma^{3}_{03}=\frac{\dot{a}}{a},$ (31d)
$\Gamma^{2}_{12}$ $=\Gamma^{3}_{13}=\frac{1}{r},\qquad\Gamma^{3}_{23}=\cot\theta,$ (31e)
$\Gamma^{1}_{00}$ $=\frac{1}{2}e^{\alpha-\beta}\alpha^{\prime},\qquad\Gamma^{1}_{11}=\frac{1}{2}\beta^{\prime},$ (31f)
$\Gamma^{1}_{22}$ $=-e^{-\beta}r,\qquad\Gamma^{1}_{33}=-e^{-\beta}r\sin^{2}\theta,$ (31g)
$\Gamma^{2}_{33}$ $=-\cos\theta\sin\theta.$ (31h)
The Einstein tensor components are
$G^{0}_{0}$ $=-\frac{e^{-\alpha}}{a^{2}}\Bigg[\frac{3\dot{a}^{2}}{a^{2}}+\frac{e^{\alpha-\beta}}{r^{2}}\Big(-1+e^{\beta}+r\beta^{\prime}\Big)+\frac{\dot{a}}{a}\dot{\beta}\Bigg],$ (32a)
$G^{0}_{1}$ $=-\frac{e^{-\alpha}}{a^{3}r}\Big(r\dot{a}\,\alpha^{\prime}+a\dot{\beta}\Big),$ (32b)
$G^{1}_{1}$ $=\frac{1}{a^{4}}\Bigg[e^{-\alpha}\dot{a}^{2}+\frac{a^{2}e^{-\beta}}{r^{2}}\Big(1-e^{\beta}+r\alpha^{\prime}\Big)+$
$+ae^{-\alpha}\Big(-2\ddot{a}+\dot{a}\,\dot{\alpha}\Big)\Bigg],$ (32c)
$G^{2}_{2}$ $=G^{3}_{3}=\frac{e^{-\alpha-\beta}}{4a^{4}r}\Bigg\{4e^{\beta}r\dot{a}^{2}+4ae^{\beta}r\dot{a}\Big(\dot{\alpha}-\dot{\beta}\Big)$
$+a\Bigg[-8e^{\beta}r\ddot{a}+a\Big[e^{\alpha}r\alpha^{\prime 2}-2e^{\alpha}\beta^{\prime}-e^{\alpha}\alpha^{\prime}(-2+r\beta^{\prime})$
$+2e^{\alpha}r\alpha^{\prime\prime}+e^{\beta}r\left(\dot{\alpha}\dot{\beta}-\dot{\beta}^{2}-2\ddot{\beta}\right)\Big]\Bigg]\Bigg\}.$ (32d)
The energy-momentum tensor is defined in Eq. (2). Using an appropriate comoving reference frame in which $u^{\mu}=(u^{0},0,0,0)$ and $w^{\mu}=(0,w^{1},0,0)$, the energy-momentum tensor components are
$T_{\mu}^{\nu}=\mbox{diag}\begin{pmatrix}-\rho,&p_{\parallel},&p_{\perp},&p_{\perp}\end{pmatrix}\,$. (33)
Using the Einstein field equations and from $T_{0}^{1}=0$ (absence of fluxes), one finds Eq. (3a). Equations (3b) and (3c) are directly derived from $G_{0}^{0}=-8\pi\rho$ and $G_{1}^{1}=8\pi p_{\parallel}$. To find the remaining equations, it is useful to consider the energy-momentum tensor conservation, which leads to the two independent equations
$\nabla^{\mu}T_{\mu 0}$ $=-\dot{\rho}-\frac{\dot{a}}{a}\left(3\rho+p_{\parallel}+2p_{\perp}\right)-\frac{\dot{\beta}}{2}(\rho+p_{\parallel})\,$, (34a)
$\nabla^{\mu}T_{\mu 1}$ $=p_{\parallel}^{\prime}+2\frac{p_{\parallel}-p_{\perp}}{r}+\frac{\alpha^{\prime}}{2}(\rho+p_{\parallel})\,.$ (34b)
Equation (3e) comes straight from Eq. (34b). At last, to find Eq. (3d), one uses Eqs. (3a) and (34b) to replace $\dot{\beta}$ from Eq. (34a).
Appendix B The Misner-Sharp mass parametrization